is c+(a+b)=(a+b)+c an example of the commutative property? is (a*b)c=a(b*c) also an example of commutative property?
Answer by HayleyAlyce(8)

c+(a+b)=(a+b)+c is an example of the associative property.
(a*b)c=a(b*c)is also an example of the associative property.